# The universal multiplicity function: counting halos and voids

### v1

@@ 3 The multiplicity function from a generic moving barrier
The exact analytical solution of the first crossing distribution exists only for a few cases, which are the ones with uncorrelated or fully-correlated steps with a constant or linear evolving threshold (Bond et al., 1991Sheth, 1998SVdWParanjape et al., 2012MS). The first crossing distribution for the general case can in principle be written as a formal expansion in an infinite series of functions involving $n-$point correlations of the field $\delta$ and its derivatives evaluated at different scales (Verechtchaguina et al., 2006; Musso & Sheth, 2014a; Nikakhtar et al., 2018), but summing this series requires approximations. Therefore, approximated methods must be used to analytically solve the Langevin equations in the excursion-set framework with correlated steps and a generic moving barrier (Peacock & Heavens, 1990; Appel & Jones, 1990; Maggiore & Riotto, 2010c; de Simone et al., 2011; Ma et al., 2011; Paranjape et al., 2012; Paranjape & Sheth, 2012; Musso et al., 2012; Musso & Sheth, 2014b, c).
@@
In the standard cosmological model, the variance of the fluctuation field, $S=\sigma^{2}(R)=C(R,R)$ (where $C(R,R)$ is a diagonal element of the covariance matrix defined in Eq. (4)), is a monotonic function of the smoothing scale $R$; therefore, these two quantities are interchangeable. To account for correlations, MS proposed a multiplicity function derived from a bivariate distribution of the density contrast filtered field and of the velocity of the filtered field with respect to the scale $S$, $\delta^{\prime}={\rm d}\delta(S)/{\rm d}S$. They replace the requirement of $\delta(s)<B(s)$ for all the scales $s<S$ (with $B(s)$ a generic moving barrier) by the milder condition $\delta(S)<B(S-\Delta S)$, with $S-s=\Delta S\rightarrow dS\rightarrow 0$. In this limit, expanding in a Taylor series both the field $\delta(S)$ and the moving barrier $B(S)$, the above condition becomes
$B_{S}\leq\delta_{S}\leq B_{S}+(\delta_{S}^{\prime}-B_{S}^{\prime})\Delta S,\quad\delta_{S}^{\prime}\geq B_{S}^{\prime},$ (7)
@@
$\begin{split}f(S)\,{\rm d}S&\simeq\int_{B^{\prime}_{S}}^{\infty}{\rm d}\delta^{\prime}\int_{B_{S}}^{B_{S}+(\delta_{S}^{\prime}-B_{S}^{\prime}){\rm d}S}{\rm d}\delta_{S}\,p(\delta_{S},\delta_{S}^{\prime})\\ &={\rm d}S\,p(B_{S})\int_{B^{\prime}_{S}}^{\infty}{\rm d}\delta^{\prime}p(\delta^{\prime}_{S}|B_{S})(\delta_{S}^{\prime}-B_{S}^{\prime})\,,\end{split}$ (8)
where $p(\delta_{S}^{\prime}|B_{S})=p(B_{S})p(B_{S},\delta_{S}^{\prime})$ is the conditional probability of $\delta_{S}^{\prime}$ at the barrier $\delta_{S}=B_{S}$ (MS). The resulting multiplicity function accurately reproduces the exact numerical solution for a general filter at large scales $R$, i.e. at small $S$, nevertheless it becomes less accurate when lowering the scale. This is expected since the correlation between different smoothing scales is modeled for an infinitesimal distance from the crossing scale, so as a local term. Spanning over a large range of scales, the local term is no longer sufficient for an accurate model, since it is not able to capture the behavior of the field at scales significantly different from the crossing.
One possibility to extend this model consists in exploring the next leading orders of the formal expansion of Musso & Sheth (2014a). Nevertheless, here we explore an alternative way which relies on accounting for correlations of the density contrast field on different scales. More precisely, we consider the joint distribution $p(\delta_{S},\delta^{\prime}_{S},\delta_{s})$, where $\delta_{s}=\delta(s)$, $S$ is the scale at which the first crossing occurs and $s<S$. In this way, by accounting for $\delta_{S}^{\prime}$, we can model the correlation on scales close to the crossing barrier, as in the MS, but accounting also for $\delta(s)$ we can model the behavior of the field at very different scales. Using the same arguments of MS, we obtain
$f(S)\,{\rm d}S\simeq\,\frac{1}{S}\int_{0}^{S}{\rm d}s\int_{-\infty}^{B(s)}{\rm d}\delta_{s}\int_{B^{\prime}_{S}}^{\infty}{\rm d}\delta^{\prime}_{S}\times$ (9)
$\qquad\qquad\left\{\int_{B_{S}}^{B_{S}+{\rm d}S(\delta^{\prime}_{S}-B^{\prime}_{S})}p(\delta_{S},\delta^{\prime}_{S},\delta_{s}){\rm d}\delta_{S}\right\}$
@@
where in the last line $\delta_{S}=B_{S}$. The integration over $\delta_{S}$, $\delta^{\prime}_{S}$, and $\delta_{s}$ represents the probability that the field $\delta$ crosses the barrier $B$ at scale $S$, without having crossed it at the larger scale $s$. Integrating over $s$, we then obtain the probability that the field $\delta$, crossing the barrier at $S$, has not crossed the barrier at any scale $s<S$. The $1/S$ term is the normalization factor: $p(\delta_{S},\delta^{\prime}_{S},\delta_{s})$ is normalized to $1$ at each $s$, therefore its integration over $s$ from 0 to $S$ gives $S$. As the field is Gaussian, the joint distribution of $\delta_{S}$, $\delta^{\prime}$, and $\delta_{s}$ is a multivariate normal distribution with covariance
${\bf\Sigma}=\begin{pmatrix}\langle\delta_{S}^{2}\rangle&\langle\delta_{S}\delta^{\prime}_{S}\rangle&\langle\delta_{S}\delta_{s}\rangle\\ \langle\delta_{S}\delta^{\prime}_{S}\rangle&\langle\delta^{\prime 2}_{S}\rangle&\langle\delta^{\prime}_{S}\delta_{s}\rangle\\ \langle\delta_{S}\delta_{s}\rangle&\langle\delta^{\prime}_{S}\delta_{s}\rangle&\langle\delta_{s}^{2}\rangle\end{pmatrix}=\begin{pmatrix}S&1/2&C\\ 1/2&D&C^{\prime}\\ C&C^{\prime}&s\end{pmatrix}.$ (10)
Here we used the fact that, by construction $\langle\delta_{S}^{2}\rangle=S$, $\langle\delta_{s}^{2}\rangle=s$, $\langle\delta_{S}\delta_{s}\rangle=C(S,s)$ defined in Eq. (4), that from now on we call simply $C$. We also define the quantities $D\equiv\langle\delta^{\prime 2}_{S}\rangle$ and $C^{\prime}\equiv\langle\delta^{\prime}_{S}\delta_{s}\rangle=\partial C/\partial S$. It can be shown that $\langle\delta_{S}\delta^{\prime}_{S}\rangle=1/2$ (see Appendix B). Concerning the quantity $\langle\delta^{\prime 2}_{S}\rangle$, in the case considered here of a top-hat filter, it is usually considered a non-convergent quantity difficult to compute due to the slowly decreasing amplitude and highly oscillating function to integrate. Nevertheless, in Appendix A we discuss the convergence of $\langle\delta^{\prime 2}_{S}\rangle$, and we show how to properly compute it. Using the properties listed in Appendix B, we obtain
@@
$\frac{1}{4\pi\sqrt{\Gamma_{\delta\delta}}}\frac{\det{\bf\Sigma}}{\Gamma_{\delta^{\prime}\delta^{\prime}}}\exp\left[-\frac{DB_{S}^{2}+SB^{\prime 2}_{S}-B^{\prime}_{S}B_{S}}{2\Gamma_{\delta\delta}}\right]\times$ (11)
$\qquad\left\{{\rm erf}\left[\sqrt{\frac{\Gamma_{\delta\delta}}{2\det{\bf\Sigma}}}\left(B(s)+\frac{\Gamma_{B\delta}}{\Gamma_{\delta\delta}}B_{S}+\frac{\Gamma_{\delta^{\prime}\delta}}{\Gamma_{\delta\delta}}B^{\prime}_{S}\right)\right]+1\right\}+$
@@
$\qquad\frac{B_{S}/2S-B^{\prime}_{S}}{4\sqrt{2\pi S}}\left[{\rm erf}\left(\sqrt{\frac{S}{2\Gamma_{\delta^{\prime}\delta^{\prime}}}}\left(B(s)-\frac{C}{S}B_{S}\right)\right)+1\right]\Bigg{\}},$
@@
$\begin{array}[]{lcl}\Gamma_{BB}=sD-C^{\prime 2}&&\Gamma_{B\delta^{\prime}}=CC^{\prime}-s/2\\ \Gamma_{\delta^{\prime}\delta^{\prime}}=sS-C^{2}&&\Gamma_{B\delta}=C^{\prime}/2-CD\\ \Gamma_{\delta\delta}=SD-1/4&&\Gamma_{\delta^{\prime}\delta}=C/2-SC^{\prime}\,,\end{array}$ (12)
@@
${\cal F}(x,\alpha,\beta)=\int_{-\infty}^{x}{\rm d}t\,t\,e^{-\alpha(t+\beta)^{2}}{\rm erf}(t)\,,$ (13)
@@
$x$ $=\frac{\Gamma_{\delta^{\prime}\delta}}{\sqrt{2\Gamma_{\delta^{\prime}\delta^{\prime}}\det{\bf\Sigma}}}\left(B(s)+\frac{\Gamma_{B\delta^{\prime}}}{\Gamma_{\delta^{\prime}\delta}}B_{S}+\frac{\Gamma_{\delta^{\prime}\delta^{\prime}}}{\Gamma_{\delta^{\prime}\delta}}B^{\prime}_{S}\right)\,,$ (14)
$\alpha$ $=\frac{\Gamma_{\delta\delta}\Gamma_{\delta^{\prime}\delta^{\prime}}}{\Gamma_{\delta^{\prime}\delta}^{2}}-1=\frac{S\det{\bf\Sigma}}{\Gamma_{\delta^{\prime}\delta}^{2}}\,,$
@@
Finally, the analytical multiplicity function is given by integrating Eq. (3): $f(S){\rm d}S={\rm d}S/S\int_{0}^{S}{\rm d}s{\cal P}(s)$. Its computation is numerically straightforward, as the quantities appearing in Eq. (3) are a combination of the elements of the covariance matrix, ${\bf\Sigma}$, which can be easily computed for any filtering function. Studying the behavior of ${\cal P}(s)$, in Appendix C we show that $\int_{0}^{S}{\cal P}(s){\rm d}s\simeq S\lim_{s\rightarrow 0}{\cal P}(s)=S\lim_{s\rightarrow S}{\cal P}(s)$. It follows that Eq. (9) can be further approximated as
$f(S)=\frac{1}{S}\int_{0}^{S}{\cal P}(s){\rm d}s\simeq$
@@
$e^{-B^{2}_{S}/2S}\left\llbracket\frac{\sqrt{\Gamma_{\delta\delta}}}{2\pi S}\exp\left[-\frac{S}{2\Gamma_{\delta\delta}}\left(\frac{B_{S}}{2S}-B^{\prime}_{S}\right)^{2}\right]\right.+$ (15)
@@
$\quad\left.\frac{\left(B_{S}/2S-B^{\prime}_{S}\right)}{2\sqrt{2\pi S}}\left\{{\rm erf}\left[\sqrt{\frac{S}{2\Gamma_{\delta\delta}}}\left(\frac{B_{S}}{2S}-B^{\prime}_{S}\right)\right]+1\right\}\right\rrbracket\,.$
The computational advantage of this approximation is remarkable, as it does not require the recomputation and integration of ${\cal P}(s)$ for each crossing scale $S$, avoiding also the numerical integration of Eq. (13), and involves only four quantities: $S$, $D$, $B_{S}$, and $B^{\prime}_{S}$.
@@
Eqs. (9)-(3)-(3) are the first of the main results presented in this work: a universal multiplicity function for a generic moving barrier, which holds both for halos and voids from small to very large scales $S$.
@@
Figure 1: Top panel: MS multiplicity function (green lines); numerical multiplicity function, Eqs. (5)–(6) (black lines); numerical integration over $s$ of Eq. (3), divided by $S$ (blue lines); corresponding analytical approximation, Eq. (3) (orange lines). Solid and dashed lines show $z=0$ and 1 results, respectively. All the multiplicity functions are obtained as a function of $S$ via the SMT at moving barrier with a top-hat filter. The upper axis shows the corresponding Lagrangian radius and mass. Middle panel: percent relative differences of the multiplicity functions in the top panel with respect to the numerical multiplicity function, Eqs. (5)–(6). The shaded area shows the accuracy of the numerical solution. Bottom panel: relative percent difference between the approximated analytical solution, Eq. (3), and the corresponding exact numerical integration over $s$ of Eq. (3), divided by $S$.
@@
The top panel of Fig. 1 shows: the MS multiplicity function (green lines); numerical multiplicity function via Monte Carlo realizations, Eqs. (5)–(6) (black lines); the exact integration of Eq. (9), i.e. the numerical integration over $s$ of Eq. (3), divided by $S$ (blue lines); the analytical approximation, given by Eq. (3), of Eq. (3) (orange lines). Solid and dashed lines show $z=0$ and 1 results, respectively. All the multiplicity functions are obtained as a function of $S$ via the SMT moving barrier with a top-hat filter. The middle panel shows the percent relative difference of the multiplicity functions in the top panel with respect to the numerical solution of Eqs. (3). The shaded area shows the accuracy of the numerical solution. The bottom panel shows the relative percent difference between the approximated analytical solution, Eq. (3), and the corresponding exact numerical integration over $s$ of Eq. (3), divided by $S$. Both the analytical solution and approximation presented here accurately reproduce the numerical solution on scales considerably smaller than the ones where the MS model is valid. In particular, at $z=0$ for halos with mass of $10^{13}~{}h^{-1}M_{\odot}$ ($5\times 10^{12}~{}h^{-1}M_{\odot}$), both our analytical solutions are accurate at $3\%$ (5%), respectively. For comparison, the MS model at the same $z$ and mass scale is accurate at $\sim 10\%$ (12.5%), respectively. Note, however, that the accuracy greatly increases with the redshift. At $z=1$ our model is precise at 1% and 2% level for masses of, respectively, $10^{13}$ and $3\times 10^{12}~{}h^{-1}M_{\odot}$, corresponding to the smallest scale appearing in Fig. 1.

### v2

@@ 3.2 A new analytical multiplicity function
The exact analytical solution of the first crossing distribution exists only for a few cases, which are the ones with uncorrelated or fully-correlated steps with a constant or linear evolving threshold [1, 40, 12, 19, 11]. The first crossing distribution for the general case can in principle be written as a formal expansion in an infinite series of functions involving $n-$point correlations of the field $\delta$ and its derivatives evaluated at different scales [41, 38, 39], but summing this series requires approximations. Therefore, approximated methods must be used to analytically solve the Langevin equations in the excursion-set framework with correlated steps and a generic moving barrier [4, 42, 43, 44, 45, 19, 16, 46, 47, 48].
@@
In the standard cosmological model, the variance of the fluctuation field, $S=\sigma^{2}(R)=C(R,R)$ (where $C(R,R)$ is a diagonal element of the covariance matrix defined in Eq. (3.3)), is a monotonic function of the smoothing scale $R$; therefore, these two quantities are interchangeable. To account for correlations, MS proposed a multiplicity function derived from a bivariate distribution of the density contrast filtered field and of the velocity of the filtered field with respect to the scale $S$, $\delta^{\prime}={\rm d}\delta(S)/{\rm d}S$. They replace the requirement of $\delta(s)<B(s)$ for all the scales $s<S$ (with $B(s)$ a generic moving barrier) by the milder condition $\delta(S)-\Delta S<B(S-\Delta S)$, with $S-s=\Delta S\rightarrow dS\rightarrow 0$. In this limit, expanding in a Taylor series both the field $\delta(S)$ and the moving barrier $B(S)$, the above condition becomes
$B_{S}\leq\delta_{S}\leq B_{S}+(\delta_{S}^{\prime}-B_{S}^{\prime})\Delta S,\quad\delta_{S}^{\prime}\geq B_{S}^{\prime},$ (3.6)
@@
$\begin{split}f(S)\,{\rm d}S&\simeq\int_{B^{\prime}_{S}}^{\infty}{\rm d}\delta^{\prime}\int_{B_{S}}^{B_{S}+(\delta_{S}^{\prime}-B_{S}^{\prime}){\rm d}S}{\rm d}\delta_{S}\,p(\delta_{S},\delta_{S}^{\prime})\\ &={\rm d}S\,p(B_{S})\int_{B^{\prime}_{S}}^{\infty}{\rm d}\delta^{\prime}p(\delta^{\prime}_{S}|B_{S})(\delta_{S}^{\prime}-B_{S}^{\prime})\,,\end{split}$ (3.7)
where $p(\delta_{S}^{\prime}|B_{S})=p(B_{S},\delta_{S}^{\prime})/p(B_{S})$ is the conditional probability of $\delta_{S}^{\prime}$ at the barrier $\delta_{S}=B_{S}$ [MS]. The resulting multiplicity function accurately reproduces the exact numerical solution for a general filter at large scales $R$, i.e. at small $S$, nevertheless it becomes less accurate when lowering the scale. This is expected since the correlation between different smoothing scales is modelled for an infinitesimal distance from the crossing scale, so as a local term. Spanning over a large range of scales, the local term is no longer sufficient for an accurate model, since it is not able to capture the behaviour of the field at scales significantly different from the crossing.
One possibility to extend this model consists in exploring the next leading orders of the formal expansion of [38]. Nevertheless, here we explore an alternative way which relies on accounting for correlations of the density contrast field on different scales. More precisely, we consider the joint distribution $p(\delta_{S},\delta^{\prime}_{S},\delta_{s})$, where $\delta_{s}=\delta(s)$, $S$ is the scale at which the first crossing occurs and $s<S$. In this way, by accounting for $\delta_{S}^{\prime}$, we can model the correlation on scales close to the crossing barrier, as in the MS, but accounting also for $\delta(s)$ we can model the behaviour of the field at very different scales. Using the same arguments of MS, we obtain
$f(S)\,{\rm d}S$ $\simeq\,\frac{1}{S}\int_{0}^{S}{\rm d}s\int_{-\infty}^{B(s)}{\rm d}\delta_{s}\int_{B^{\prime}_{S}}^{\infty}{\rm d}\delta^{\prime}_{S}\int_{B_{S}}^{B_{S}+{\rm d}S(\delta^{\prime}_{S}-B^{\prime}_{S})}p(\delta_{S},\delta^{\prime}_{S},\delta_{s}){\rm d}\delta_{S}$ (3.8)
@@
where in the last line $\delta_{S}=B_{S}$. The integration over $\delta_{S}$, $\delta^{\prime}_{S}$, and $\delta_{s}$ represents the probability that the field $\delta$ crosses the barrier $B$ at scale $S$, without having crossed it at the larger scale $s$. Integrating over $s$, we then obtain the probability that the field $\delta$, crossing the barrier at $S$, has not crossed the barrier at any scale $s<S$. The $1/S$ term is the normalisation factor: $p(\delta_{S},\delta^{\prime}_{S},\delta_{s})$ is normalised to $1$ at each $s$, therefore its integration over $s$ from 0 to $S$ gives $S$. As the field is Gaussian, the joint distribution of $\delta_{S}$, $\delta^{\prime}$, and $\delta_{s}$ is a multivariate normal distribution with covariance
${\bf\Sigma}=\begin{pmatrix}\langle\delta_{S}^{2}\rangle&\langle\delta_{S}\delta^{\prime}_{S}\rangle&\langle\delta_{S}\delta_{s}\rangle\\ \langle\delta_{S}\delta^{\prime}_{S}\rangle&\langle\delta^{\prime 2}_{S}\rangle&\langle\delta^{\prime}_{S}\delta_{s}\rangle\\ \langle\delta_{S}\delta_{s}\rangle&\langle\delta^{\prime}_{S}\delta_{s}\rangle&\langle\delta_{s}^{2}\rangle\end{pmatrix}=\begin{pmatrix}S&1/2&C\\ 1/2&D&C^{\prime}\\ C&C^{\prime}&s\end{pmatrix}.$ (3.9)
Here we used the fact that, by construction $\langle\delta_{S}^{2}\rangle=S$, $\langle\delta_{s}^{2}\rangle=s$, $\langle\delta_{S}\delta_{s}\rangle=C(S,s)$ defined in Eq. (3.3), that from now on we call simply $C$. We also define the quantities $D\equiv\langle\delta^{\prime 2}_{S}\rangle$ and $C^{\prime}\equiv\langle\delta^{\prime}_{S}\delta_{s}\rangle=\partial C/\partial S$. It can be shown that $\langle\delta_{S}\delta^{\prime}_{S}\rangle=1/2$ (see Appendix B). Differently from previous works, here we evaluate exactly the quantity $\langle\delta^{\prime 2}_{S}\rangle$. In fact, for a of a top-hat filter, this quantity is usually considered non-convergent, due to the slowly decreasing amplitude and highly oscillating functions to integrate. In Appendix A we discuss the convergence of $\langle\delta^{\prime 2}_{S}\rangle$, showing how to properly compute it. Using the properties listed in Appendix B, we obtain
@@
$\frac{1}{4\pi\sqrt{\Gamma_{\delta\delta}}}\frac{\det{\bf\Sigma}}{\Gamma_{\delta^{\prime}\delta^{\prime}}}\exp\left[-\frac{DB_{S}^{2}+SB^{\prime 2}_{S}-B^{\prime}_{S}B_{S}}{2\Gamma_{\delta\delta}}\right]\times$ (3.10)
$\qquad\qquad\left\{{\rm erf}\left[\sqrt{\frac{\Gamma_{\delta\delta}}{2\det{\bf\Sigma}}}\left(B(s)+\frac{\Gamma_{B\delta}}{\Gamma_{\delta\delta}}B_{S}+\frac{\Gamma_{\delta^{\prime}\delta}}{\Gamma_{\delta\delta}}B^{\prime}_{S}\right)\right]+1\right\}+$
@@
$\qquad\qquad\frac{B_{S}/2S-B^{\prime}_{S}}{4\sqrt{2\pi S}}\left[{\rm erf}\left(\sqrt{\frac{S}{2\Gamma_{\delta^{\prime}\delta^{\prime}}}}\left(B(s)-\frac{C}{S}B_{S}\right)\right)+1\right]\Bigg{\}},$
@@
$\begin{array}[]{lcl}\Gamma_{BB}=sD-C^{\prime 2}&&\Gamma_{B\delta^{\prime}}=CC^{\prime}-s/2\\ \Gamma_{\delta^{\prime}\delta^{\prime}}=sS-C^{2}&&\Gamma_{B\delta}=C^{\prime}/2-CD\\ \Gamma_{\delta\delta}=SD-1/4&&\Gamma_{\delta^{\prime}\delta}=C/2-SC^{\prime}\,,\end{array}$ (3.11)
@@
${\cal F}(x,\alpha,\beta)=\int_{-\infty}^{x}{\rm d}t\,t\,e^{-\alpha(t+\beta)^{2}}{\rm erf}(t)\,,$ (3.12)
@@
$x$ $=\frac{\Gamma_{\delta^{\prime}\delta}}{\sqrt{2\Gamma_{\delta^{\prime}\delta^{\prime}}\det{\bf\Sigma}}}\left(B(s)+\frac{\Gamma_{B\delta^{\prime}}}{\Gamma_{\delta^{\prime}\delta}}B_{S}+\frac{\Gamma_{\delta^{\prime}\delta^{\prime}}}{\Gamma_{\delta^{\prime}\delta}}B^{\prime}_{S}\right)\,,$
$\alpha$ $=\frac{\Gamma_{\delta\delta}\Gamma_{\delta^{\prime}\delta^{\prime}}}{\Gamma_{\delta^{\prime}\delta}^{2}}-1=\frac{S\det{\bf\Sigma}}{\Gamma_{\delta^{\prime}\delta}^{2}}\,,$ (3.13)
@@
Finally, the analytical multiplicity function is given by integrating Eq. (3.2):
$f(S)\,{\rm d}S=\frac{{\rm d}S}{S}\int_{0}^{S}{\rm d}s\,{\cal P}(s)\,.$ (3.14)
@@
Its computation is numerically straightforward, as the quantities appearing in Eq. (3.2) are a combination of the elements of the covariance matrix, ${\bf\Sigma}$, which can be easily computed for any filtering function. Studying the behaviour of ${\cal P}(s)$, in Appendix C we show that $\lim_{s\rightarrow 0}{\cal P}(s)=\lim_{s\rightarrow S}{\cal P}(s)$. It follows that, in the small $S$ limit, Eq. (3.14) can be further approximated as
@@
$f(S)\simeq\frac{e^{-B^{2}_{S}/2S}}{\sqrt{2\pi S}}$ $\left\llbracket\sqrt{\frac{\Gamma_{\delta\delta}}{2\pi S}}\exp\left[-\frac{S}{2\Gamma_{\delta\delta}}\left(\frac{B_{S}}{2S}-B^{\prime}_{S}\right)^{2}\right]\right.+$
@@
$\ \left.\frac{1}{2}\left(\frac{B_{S}}{2S}-B^{\prime}_{S}\right)\left\{{\rm erf}\left[\sqrt{\frac{S}{2\Gamma_{\delta\delta}}}\left(\frac{B_{S}}{2S}-B^{\prime}_{S}\right)\right]+1\right\}\right\rrbracket\,.$ (3.15)
This equation is formally the MS multiplycity function, with the difference that in this work the $\Gamma_{\delta\delta}$ term, which depends on $D$, is exactly computed accounting for its scale dependence rather than using a fixed approximated value. This approximation is quite accurate at medium and large scales, involving only four quantities: $S$, $D$, $B_{S}$, and $B^{\prime}_{S}$.
@@
Eqs. (3.2)-(3.14)-(3.2) are the first of the main results presented in this work: a universal multiplicity function for a generic moving barrier, which holds both for haloes and voids from small to very large scales $S$. In Appendix D we discuss the local bias expansion in the effective barrier approach.
@@
Figure 1: Top panel: numerical multiplicity function via MC realisations, Eqs. (3.4)–(3.5) (black lines); analytical multiplicity function, Eq. (3.14) (blue lines); corresponding approximation, Eq. (3.2) (orange lines); the MS multiplicity function (see text, green lines). Solid and dashed lines show $z=0$ and 1 results, respectively. All the multiplicity functions are obtained as a function of $S$ via the SMT moving barrier with the top-hat filter. The upper axis shows the corresponding Lagrangian radius and mass. Middle panel: percent relative differences of the multiplicity functions in the top panel with respect to the numerical multiplicity function, Eqs. (3.4)–(3.5). The shaded and hatched area show the accuracy of the numerical solution. Bottom panel: relative difference between the analytical multiplicity functions, Eqs. (3.14)–(3.2) and MS, and the numerical one in $S$ units.
@@
The top panel of Fig. 1 shows: the numerical solution of Eqs. (3.2) via Monte Carlo (MC) realisations, Eqs. (3.4)–(3.5) (black lines); the exact integration of Eq. (3.8), i.e. Eq. (3.14) (blue lines); the corresponding approximation, given by Eq. (3.2) (orange lines); the MS multiplicity function, with $\Gamma$ appearing in their Eq. (5) fixed to $1/3$ [MS], which correspond to $\Gamma_{\delta\delta}=3/4$ in Eq. (3.2) (green lines). Solid and dashed lines show $z=0$ and 1 results, respectively. All the multiplicity functions are obtained as a function of $S$ via the SMT moving barrier with the top-hat filter. The middle panel shows the percent relative difference of the multiplicity functions in the top panel with respect to the numerical solution. The shaded and hatched areas show the accuracy of the numerical solution for the number of MC realisations used, at $z=0$ and 1 respectively. To better visualise the deviation from the numerical solution, the bottom panel shows the relative difference between the analytical multiplicity functions, Eqs. (3.14)–(3.2) and MS, and the numerical one in $S$ units, i.e. the relative difference divided by $S$. Note the importance of considering the exact $D$ scale dependence: both Eq. (3.14) (blue) and Eq. (3.2) (orange) accurately reproduce the numerical solution on scales considerably smaller than the MS model with $\Gamma_{\delta\delta}$ fixed at $3/4$ (green). In particular, at $z=0$ for haloes with mass of $10^{14}$, $10^{13}$, and $5\times 10^{12}~{}h^{-1}M_{\odot}$, Eq. (3.14) is accurate at $\sim 1\%$, 3% and 5%, while Eq. (3.2) is accurate at $\sim 1.5\%$, 5%, and 7%, respectively, i.e. roughly $50\%$ less accurate than Eq. (3.14). For comparison, the MS model ($\Gamma_{\delta\delta}$ fixed) is accurate at $\sim 5\%$, 8.5%, and 9%, at the same mass scales. Note, however, that the accuracy largely increases with the redshift: at $z=1$ for the same mass scales Eq. (3.14) is accurate at $\sim 0.25\%$, 1% and 1.5%, while Eq. (3.2) is accurate at $\sim 0.25\%$, 1.5%, and 2%, respectively.